what is the wavelength of a proton(mass =1.673 x 10e-27kg) traveling at a speed of 4.55 x 10e6 m/s?
1
Expert's answer
2020-10-12T13:35:45-0400

Given mass of proton, M = 1.673 x 10e-27kg

speed of proton, V = 4.55 x 10e6 m/s



Momentum of this proton, p = MV kg.m/s

= 1.673 x 10e-27 × 4.55×10e6 kg.m/s

=7.612 × 10e-21 kg.m/s



the wavelength of a proton, λ = (h/p) unit

(Where, h = plank's constan )

Putting h = 6.626 × 10e-34 J.s

and p = 7.612 × 10e-21 kg.m/s

λ = h/p =(6.626 × 10e-34) / (7.612 × 10e-21) m

= 8.704 × 10e-14 m


Hence, the wavelength of a proton is = 8.704 × 10e-14 m